Carbon temachlonde reacts at high temperatures wth oygento produce two toxic geses phosgene and chlorine. What K for the following reaction? O 6.6 x 10 O 44 10 O 88 x 109 O 19 x 10 2.3 x 10-10
Answer #1

New Kc "= Kc ^2

= 4.4 * 10^9) ^ 2

= 1.9 * 10^ 19 ( option D )

Since new reaction is double so equibirum constant will be square of constant.